answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the PROBABILITY of throwing 2 sixes with two dice?

This is an example of an INDEPENDENT event:

The probability of throwing a double six with two dice is the result of throwing six with the first die and six with the second die. The total possibilities are, one from six outcomes for the first event and one from six outcomes for the second, Therefore (1/6) * (1/6) = 1/36th or 2.77%.

The two events are independent, since whatever happens to the first die cannot affect the throw of the second, the probabilities are therefore multiplied, and remain 1/36th.

What is the sample space of tossing 4 coins?

If the order of the outcomes matters, then

TTTT,

TTTH, TTHT, THTT, HTTT,

TTHH, THTH, THHT, HTTH, HTHT, HHTT,

THHH, HTHH, HHTH, HHHT,

HHHH.

If the order does not matter, then

TTTT, TTTH, TTHH, THHH AND HHHH

How do you solve problems on probability?

There are many different problems and different ways for solving them.

3 cards are drawn simultaneously at random from a standard deck Find the probability that all the cards are hearts?

Here number of sample points that is total possible outcomes would be 52C3.
Event A = Three cards draw are hearts.
So here we can select 3 hearts out of 13 hearts in 13C3 ways...

Therefore the required probability would be
=(13C3)/(52C3)
=286/22100
=0.01294

Kunal K.

14 players for a softball team show up for a game How many ways are there to assign the 10 positions by selecting players from the 14 people who show up?

Try calling it, "14 choose 10". Set up the equation: n!/r!(n-r)!

n=14, r=10 So: (14!)/[10!(14-10)!]

Final answer: 1001 different ways. The number is so large it is simply counter intuitive, isn't it?....but it's true!

Probability of two tetrahedral dice rolled the sum on the upward faces is four?

When a tetrahedral die is rolled, it will rest with three faces upwards. If the die is numberd from 1 to 4. therefore the sum of the upward facing numbers on 1 die is at least 6 and so for two dice, the minimum is 12. That being the case, the probability is 0.

When a single card is drawn from an ordinary 52 card deck what is the probability of drawing a 3 or a club?

In an ordinary 52 card deck, there are 4 three's, and 13 clubs, one of which is the 3 of clubs, so there are 16 cards that are a three or a club. The probability of drawing a three or a club is 16 out of 52, or 4 out or 13, or 0.308.

If 4 coins are tossed what is the probability of getting 3 heads?

C1, C2, C3, C4


The number of combinations is 4 base 2, or 16.


There are 4 possible scenarios where you get 3 heads.




So the odds are 4:16, or 1:4.

If five dice are rolled how many possible outcomes?

Depends on what detail of outcome you are interested in. If you are only interested in the total, you can get all numbers between 5 (5 ones) and 30 (5 sixes) - a total of 26 different results.

Probability of having 3 girls out of 4 children?

There is no simple answer to the question because the children's genders are not independent events. They depend on the parents' ages and their genes.

However, if you assume that they are independent events then, given that the probability of a boy is approx 0.52, the probability of 3 girls out of 4 children is 0.2331

A bag that has 7 greens marbles 3 blue marbles and 12 white marbles what is the probability of selecting a white marble on each of the three tries if the marble is placed back after each pick?

12 white marbles from (7+3+12) = 22 marbles

Chance of a white marble on first pick = 12/22 = 6/11.
Chance of a white marble on second and third picks are the same, as the marble is replaced.
So, the chance of a white marble three times is 6/11 * 6/11 * 6/11 = 216/1331 = approximately 16.23%

How do calculate the probability of at most?

For a discrete variable, you add together the probabilities of all values of the random variable less than or equal to the specified number.

For a continuous variable it the integral of the probability distribution function up to the specified value.

Often these values may be calculated or tabulated as cumulative probability distributions.

Trending Questions
What is the probability of choosing a month starting with the letter m? What is the result of dividing the number of successful outcomes by the number of possible outcomes? What is the probability of rolling a 6 and a 4? Combinatios of 5 in 36 numbers? What is the meaning of Sampling distribution of the test statistic? What is the probability that at least 2 people have the same birth month in a group of 8 people? In how many different orders can 4 people be arranged? Is the area under the standard normal distribution to the left of z equals 0 negative? If a person is randomly selected from the US population the odds the person lives in California are 1 to 8 What is the probability to two decimal places of a randomly chosen person being from Cali? The number of ways 5 cards can be dealt from a standard deck of cards? Probability distribution of four coins being tossed? Kinds of non random sampling technique? What we do if 0 percent basophil? What is the probability that exactly four vehicles will pass the nct test if a recent study reveals that 80 percent of all vehicles passed the nct test a random sample of 6 vehicles is selected? After 5 points are added to every score in a distribution the mean is calculated and found to be m equals 30 What was the value of the mean for the original distribution? Why is only one normal table need to find any probability under the normal curve? What is the probability of rolling a number less than 3 and greater than 4 with one roll of a number cube? How many times does head and A tail if you flip a coin? Why is standard deviation of a statistic called standard error? The probability of an impossible event is ______ or ____.?